### v1

# Direct-to-Cell: A First Look into Starlink’s Direct Satellite-to-Device Radio Access Network through Crowdsourced Measurements

Abstract
Low Earth Orbit (LEO) satellite mega-constellations have recently emerged as a viable access solution for broadband services in underserved areas. In 2024, Direct Satellite-to-Device (DS2D) communications, which enable unmodified smartphones to connect directly to spaceborne base stations, entered large-scale beta testing, with Starlink globally leading deployments. This paper presents the first measurement study of commercial DS2D services. Using crowdsourced mobile network data collected in the U.S. between October 2024 and April 2025, our research derives evidence-based insights into the capabilities, limitations, and prospective evolution of DS2D technologies providing Supplemental Coverage from Space (SCS) services to expand existing mobile network connectivity. We observe a strong correlation between the number of satellites deployed and the expanding extension of observed measurements, concentrated in accessible but poorly covered areas by terrestrial networks, such as national parks and large low-density counties. The data reveal stable physical-layer value measurement throughout the observation period, with a lower median RSRP (24-dB difference) and a higher RSRQ (3-dB difference) compared to terrestrial networks, reflecting the SMS-only usage of the DS2D network during this period. Based on SINR measurements, we estimate the expected performance of the announced DS2D mobile data service to be around 4 Mbps per beam in outdoor conditions. We also discuss strategies to expand this capacity up to 24 Mbps in the future, depending on key regulatory decisions regarding satellite licenses, spectrum availability, and allowable radiated power levels.
Index Terms:
Direct Satellite-to-Device, Starlink, Measurements, Crowdsourced, LEO, LTE, cellular networks, NTN
I Introduction
Over the past decade, mobile networks have become an essential means of accessing the Internet for billions of users worldwide. Despite a remarkable growth of mobile network coverage, significant portions of the global population remain underserved by terrestrial cellular infrastructure, particularly in remote, rural, and sparsely populated areas where the cost of deployment is prohibitive.
In recent years, particularly from 2020, Low Earth Orbit (LEO) mega-constellations have become a viable and increasingly widespread Non-Terrestrial Network (NTN) internet access solution for fixed broadband services. Networks like SpaceX’s Starlink, Eutelsat’s OneWeb or Amazon’s Kuiper have successfully addressed longstanding limitations of traditional geostationary satellite communications. By operating at significantly lower altitudes (typically 5005005005001,20012001,2001 , 200 km), these constellations offer latency levels comparable to terrestrial networks (typically tens of milliseconds vs. the hundreds associated with geostationary satellites), allowing real-time applications such as video conferencing and online gaming [1]. Furthermore, the growing number of satellites in orbit thanks to the plummeting deployment costscombined with the use of advanced beamforming technologies has contributed to significantly increasing the capacity of the system through densification and enhanced spectrum reuse.
Following these developments in fixed broadband connectivity, satellite-based cellular communications, known as Direct Satellite-to-Device (DS2D), have recently emerged as a promising solution to extend mobile network coverage. Taking advantage of the technological advancements demonstrated by LEO constellations [2], these systems aim to provide seamless connectivity directly to standard unmodified smartphones. Starlink, the satellite communications division of SpaceX, has become the global leader in DS2D communications. In partnership with T-Mobile, it has been the first to launch large-scale beta testing of a Supplemental Coverage from Space (SCS) service using T-Mobile’s spectrum, initially limited to short messages (SMS) but planned to support voice and data services by mid-2025.
In this article, we examine Starlink’s SCS services and characterize its DS2D network —–commonly referred to by Starlink as Direct-to-Cell. Specifically, we evaluate its use by observing metrics of LTE (4G) access networks based on physical-layer measurements, namely Reference Signal Received Power (RSRP), Reference Signal Received Quality (RSRQ) and Signal-to-Interference-plus-Noise Ratio (SINR). To this end, we leverage a large crowdsourced mobile network measurement dataset provided by Weplan Analytics, comprising more than one million measurements of Starlink’s DS2D network collected across the U.S. between October 2024 and April 2025. Our analysis offers valuable insights into the implications of DS2D network design for mobile network planning and spectrum policy and allows projecting the performance of future data services.
The remainder of this paper is structured as follows. Section II provides background on the evolution of LEO satellite networks and summarizes related research. Section III describes the dataset and our analytical approach. Section IV presents and discusses the results. Finally, Section V offers conclusions and suggests future research.
II Background
II-A Evolution of LEO satellite Communication Networks
The evolution of LEO satellite communications has progressed considerably in recent decades, driven by satellite miniaturization, improved launch capabilities, and the growing demand for ubiquitous connectivity. Early LEO constellations in the late 1990s were primarily designed to deliver voice and low-speed data services to remote and underserved areas. These systems demonstrated the feasibility of low-latency real-time communication via LEO satellites; however, their high deployment and operational costs and limited capacity hindered large-scale commercial adoption.
A new phase in LEO communications emerged in the late 2010s with the advent of mega-constellations. Initiatives such as SpaceX’s Starlink and OneWeb began deploying extensive LEO satellite networks to provide broadband internet access globally. In particular, Starlink has significantly transformed the sector by launching thousands of satellites in a short period of time. The first ones were deployed in 2019, and, by the end of 2024, the company operated more than $7,000$. Initially, Starlink adopted a bent-pipe architecture [3], with satellites acting as relays that amplify and forward user signals to ground stations without demodulating them. This approach enabled rapid deployment using simpler satellite designs that accelerated rollouts. However, this model imposed notable constraints, including the need for a dense network of ground stations to ensure continuous coverage and the inability to serve areas where no ground station was visible to the satellite, such as over oceans and isolated regions.
To overcome these limitations, Starlink initiated the integration of laser communications into its later first-generation satellites from 2022, and fully incorporated this technology into its second-generation (Gen2) fleet to enable Inter-Satellite Links (ISLs). This architectural advancement allowed satellites to demodulate signals onboard and dynamically route data between satellites in orbit, alleviating the need for visibility of a ground station. With this, Starlink has evolved from a ground-reliant relay network to a hybrid space-based mesh architecture, expanded its geographical reach, and improved the resilience and end-to-end performance of its network[1].
II-B Direct Satellite-to-Device (DS2D)
Building on general advances in LEO satellite communications, DS2D has been further enabled by significant progress in electronics, embedded systems, digital signal processing, and beamforming technologies, among others [4]. Specifically, DS2D connectivity relies on the satellites being equipped with advanced phased array antennas, highly sensitive radio receivers, and high-power transmitters. Moreover, compensation for Doppler shift and delay is essential to overcome the physical constraints of mobile-to-satellite links, and software-defined radio (SDR) is necessary to adapt the system to varying frequencies and technologies. Starlink’s DS2D architecture leverages an LTE eNodeB payload on board, enabling direct connections from smartphones using terrestrial mobile standards. Traffic can be routed either through the company’s existing network of ground stations or, when necessary, across the constellation via ISLs [5].
SpaceX announced the completion of its first DS2D constellation in December 2024, consisting of approximately 400 satellites –barely 6% of their deployed fleet. In addition to the SCS service provided in the U.S. through the partnership with T-Mobile, the constellation also offers SCS connectivity in New Zealand via a partnership with One NZ and is currently under testing with operators in Australia, Canada, Chile, Peru, and Japan.
In the U.S., the FCC authorized Starlink to use a $2\times 5$ MHz channel in the 1900 MHz PCS G Block (19101910191019101915191519151915 MHz uplink, 19901990199019901995199519951995 MHz downlink) for DS2D communications, subject to restrictions that prevent interference with primary terrestrial mobile services. [^1] See FCC’s IBFS File No.: SAT-LOA-20210511-00064 This spectrum is licensed to T-Mobile across the Contiguous U.S.
In November 2024, the FCC granted partial approval for Starlink’s DS2D operations to initiate the SMS-only beta testing but imposed several conditions to prevent harmful interference with terrestrial and satellite incumbents, including the requirement to stop transmission immediately upon any substantiated claim of interference, limitations on power flux density levels for aggregate out-of-band emissions (OOBE), and mandatory coordination procedures with incumbent licensees. [^2] See FCC’s DA 24-1193 Full approval for broader services, such as voice and data, was deferred pending further demonstration of interference mitigation.
Subsequently, in March 2025, the FCC approved a conditional waiver allowing Starlink to increase its OOBE by $10$ dB [^3] See FCC’s DA 25-197, despite formal objections from terrestrial operators, including AT&T and Verizon, and from EchoStar. The waiver stipulated that Starlink must continue to adhere to strict OOBE limits and engage in active interference resolution with affected parties.
II-C Related Work
Mirroring the rapid pace of innovation of LEO satellite communication technologies, a growing body of research has emerged from multiple perspectives to explore the design, deployment, and operation of these systems. Early studies focused primarily on the foundational technologies that underpin LEO broadband services, addressing alternative network architectures (e.g., bent-pipe versus ISL designs), hybrid beamforming strategies, interference coordination mechanisms, and resource management challenges [2]. Although much of this work was initially centered on fixed satellite broadband services, recent attention has shifted toward DS2D communications [6], which is reflected in dedicated research and publication initiatives (see e.g. [7]). Alongside theoretical and architectural advances, new simulation tools, such as Hypatia and StarPerf, have been developed to support realistic modeling, design, and performance evaluation of LEO satellite networks under operational conditions.
Following initial deployments and the establishment of operational networks, a growing number of measurement studies has emerged, primarily focusing on Starlink’s fixed satellite broadband services (see e.g. [1, 8, 9, 10]). These studies have typically evaluated connectivity delivered through dedicated fixed user terminals (i.e., satellite dishes), with particular attention to metrics such as throughput and, notably, latency, a major historical limitation of traditional geostationary satellite systems.
Despite the growing body of literature on LEO satellite networks, there are, to the best of our knowledge, no empirical evaluations of commercial DS2D deployments. This paper presents the first measurement-based analysis of a real-world DS2D network, leveraging large-scale crowdsourced mobile data collected during the beta testing phase of Starlink’s service in the U.S.
III Methodology
We adopt a data-driven, user-centric methodology in this study, and leverage large-scale crowdsourced mobile measurements provided by the crowdsourcing company Weplan Analytics, whose data collection methodology is specifically optimized for evaluating Radio Access Network (RAN) performance. Crowdsourced measurements offer a unique vantage point for evaluating cellular network performance, capturing real-world user experiences even in sparsely populated areas [11]. Subsection III-A describes the data sources and the data preprocessing; Subsection III-B presents how we estimate the use of SCS services; and Subsection III-C details how we calculate the expected performance of DS2D communications based on the SINR measurements.
III-A Data sources
Our primary data source consists of LTE crowdsourced measurements passively collected from Android user devices between October 2024 and April 2025. All data were anonymized at collection time by the crowdsourcing provider to ensure full compliance with applicable data protection regulations. For this study, we used only the following network variables, most of them defined by the 3GPP specifications [12].
  • PLMN (Public Land Mobile Network): identifier of the mobile network through the Mobile Country Code (MCC) and Mobile Network Code (MNC). We filter with $\mathrm{MCC}=310$ and $\mathrm{MNC}=830$ or $210$ for Starlink measurements, and $\mathrm{MNC}=260$ for T-Mobile’s terrestrial network, which will be used as a reference to benchmark Starlink’s metrics with those of terrestrial systems.
  • EARFCN (E-UTRA Absolute Radio Frequency Channel Number).
  • Cell bandwidth: $5$ MHz in the case of the PCS G Block.
  • ECI (E-UTRAN Cell Identifier).
  • RSRP: power received from the Cell-specific Reference Signals (CRS).
  • RSRQ: quality indicator computed as $N\cdot\mathrm{RSRP}/\mathrm{RSSI}$, where RSSI (Received Signal Strength Indicator) is the total power (including interference and noise) received within the cell’s bandwidth, and $N$ is the corresponding number of resource blocks.
  • SINR: unlike RSRP and RSRQ, SINR is only measured by the terminal when there is an active connection.
  • Timestamp: Date and time of each measurement.
  • GPS location: Latitude and longitude of the user terminal.
In addition, we incorporate three complementary data sources that help assess the consistency and reliability of the crowdsourced dataset. First, we compiled a dataset of the cumulative number of DS2D-capable satellite launches per date by aggregating information from official SpaceX mission reports, which we publish openly as supplementary material of this paper. This allows us to compare the observed growth of the DS2D network, reflected in the increasing number of unique measured ECIs in the crowdsourced data, with the actual deployment timeline of satellite launches. Second, we use the publicly available coverage map from T-Mobile (as of April 2025) to evaluate how well the spatial distribution of DS2D measurements aligns with areas designated as lacking terrestrial coverage, and to assess the effective coverage enhancement added by the SCS service. Finally, we used official county-level population data from the U.S. Census Bureau to quantify the intensity of SCS service usage, as detailed next.
III-B Measuring SCS use
To quantify the relative use of SCS services across the U.S., we compute the ratio of Starlink DS2D measurements to the total number of measurements, defined as the sum of Starlink DS2D and T-Mobile terrestrial network measurements across all technologies and frequency bands. This ratio is calculated at the county level, providing a balance between geographic granularity and statistical robustness, helping to mitigate potential micro-level biases inherent in crowdsourcing methodologies. The resulting indicator ranges from 00 (counties with no observed Starlink measurements) to $1$ (counties with no observed T-Mobile terrestrial measurements), effectively capturing the reliance on SCS in different geographic areas.
III-C Estimating DS2D performance
[…]
IV Results
This section presents our empirical findings. We start by assessing the relationship between the observed number of cell identifiers and the satellite launches in Subsection IV-A. Susbsection IV-B then examines the spatial availability of Starlink’s SCS service, characterizing the evolution of its geographic footprint, and quantifying use across different types of areas. Susbsection IV-C analyzes key RAN performance indicators, including RSRP, RSRQ, and SINR. Finally, we estimate the expected capacity of DS2D data services under current conditions based on these metrics, and discuss and quantify alternatives to improve it in Susbsection IV-D.
IV-A Network Expansion vs. Constellation Deployment
Figure 1: Monthly unique ECI vs. nº operational satellites.
Fig. 1 shows the evolution of the number of cumulative unique cell identifiers (ECIs) associated with Starlink’s DS2D network, as observed in our dataset over the measurement period. The results reveal a steady upward trend, closely correlated with the timeline of Gen2 satellite launches. Next we study the spatio-temporal distribution of these measurements.
This finding suggests that crowdsourcing may serve as a useful tool for large-scale characterization of DS2D networks and challenges the common perception that such data are inherently biased toward urban areas. On the contrary, these results indicate that crowdsourced measurements can offer valuable insights even in sparsely populated regions, where SCS services are likely to have the greatest impact.
IV-B Starlink’s DS2D Spatial Analysis
Although satellite constellations are designed to provide global coverage by definition, the actual use of SCS services is not geographically uniform. In practice, service activation and usage patterns vary significantly across regions, influenced by factors such as terrestrial network availability, accessibility, regulatory constraints, and operator deployment strategies.
Figure 2: Spatio-temporal evolution of observed Starlink’s DS2D measurements from October 2024 to April 2025.
Fig. 2 shows the evolution of the spatial distribution of Starlink’s DS2D mobile crowdsourced measurements, highlighting the progressive activation of the service across the United States. During the fall of 2024 (left), significant activity was recorded in Florida, Georgia, and North and South Carolina, after the activation of emergency services authorized by the FCC in response to hurricanes Helene and Milton. In the aftermath of these storms, terrestrial telecoms infrastructure in many areas became inoperative, and Starlink’s DS2D service provided temporary mobile connectivity free of charge to all mobile users, regardless of their network provider. A similar pattern emerged during the January 2025 Southern California wildfires, located in Los Angeles county and neighboring ones.
The measurements of early 2025 (center), which include the official beta service launch in January 2025, show an expanded geographic distribution nationwide. Finally, the data corresponding to March–April 2025 (right) display a similar spatial footprint to that observed earlier in the year, although with a significantly higher number of measurements, underscoring increased adoption following the beta launch. Moreover, the results indicate that service activation was geographically controlled, likely through geofencing and beamforming, during early deployment phases. Prior to the constellation’s completion in December 2024, the data show that connectivity was selectively enabled in specific regions.
In Fig. 3, we present a logarithmic-scale heatmap showing the number of observed Starlink DS2D measurements overlaid on T-Mobile’s public coverage map. The comparison reveals noteworthy insights. First, DS2D measurements occur primarily in areas designated by T-Mobile as having satellite-only coverage. Interestingly, some measurements also appear in regions officially covered by terrestrial networks, which points to poor or unreliable performance or terrestrial networks; however, analyzing this aspect lies beyond the scope of the present study and will be addressed in future works.
Figure 3: Starlink’s measurement counts vs. T-Mobile’s coverage.
Second, while DS2D is primarily used in areas with low overall population density, most measurements are concentrated in accessible yet underserved regions near populated zones. This is particularly evident around national parks in the western U.S., such as Mendocino National Forest, Redwood National and State Parks, and Rocky Mountain National Park, where we qualitatively observe significant DS2D usage. In addition, we conduct a quantitative analysis using the SCS usage indicator introduced in Section III-B to measure the relative intensity of Starlink’s DS2D service. We find a moderately strong negative logarithmic correlation with county-level population density (Pearson coefficient: $-0.59$), as shown in Fig. 4. Counties such as Knott, KY; Florence, WI; Hancock, TN; and Martin, KY, also located near natural parks, exhibit usage intensities exceeding 85%.
Figure 4: Intensity of SCS use vs. population density by county.
IV-C Analysis of DS2D Key Performance Indicators
In this subsection, we present key metrics of Starlink’s DS2D RAN, namely RSRP, RSRQ and SINR, that offer a hint on the expected performance of future Starlink’s DS2D data services.
Fig. 5 shows biweekly distributions (in the form of letter-value plots) of RSRP, RSRQ and SINR measurements for Starlink’s DS2D network over time. The chart spans the different phases of the beta testing period, with each bar representing two weeks of data, from the first week of October 2024 to the last week of April 2025. This timeline captures key milestones, including the initial activation of emergency services in hurricane-affected areas in early October, the official launch of the beta testing in January 2025, and the authorization granted in March 2025 to increase OOBEs.
Figure 5: Biweekly letter-value plots of key RAN metrics.
Figure 6: CDF of physical-layer measurements for Starlink DSD and T-Mobile’s terrestrial networks.
However, a detailed analysis of these metrics reveals no significant variation or identifiable trend in the distribution of RSRP, RSRQ, or SINR measurements over the analyzed period. On the contrary, the DS2D RAN metrics remained relatively stable throughout the different deployment phases. Nonetheless, these changes are not reflected in the measured RAN indicators, likely because the network remained largely unloaded, as services were limited to SMS during this period.
A key question is how this DS2D RAN behavior differs from that based on terrestrial cellular technologies. For this, we benchmark the metrics of this network with those of T-Mobile’s LTE terrestrial network. Fig. 6 shows the cumulative distribution functions (CDF) of RSRP, RSRQ and SINR observed in Starlink DS2D and in the terrestrial network, over the full time span of the measurements. As can be seen, users of the Starlink DS2D network experience significantly lower RSRP levels, with median values around $-121$ dBm, compared to the approximately $-97$ dBm measured in T-Mobile’s terrestrial network. This disparity is primarily due to the much larger transmission distances in satellite communications, which result in DS2D RSRP values that resemble those typically observed at the cell edge in terrestrial networks. The distribution is also narrower: the interdecile range of RSRP measurements in the DS2D network is $21$ dB, vs. $37$ dB in terrestrial networks. This reflects the more uniform link geometry: unlike terrestrial networks —where users may be located at highly varying distances from base stations and subject to diverse propagation conditions, including outdoor and indoor— DS2D users receive signals from satellites at more consistent distances and predominantly under line-of-sight conditions.
Regarding RSRQ, despite the limited bandwidth available to the DS2D network ($5$ MHz), values are generally more favorable for DS2D connections (median: $-9$ dB) than in terrestrial networks ($-12$ dB). This is explained by the low network load and minimal interference from other active terminals within the same coverage area, due to the current usage of the DS2D network: operation in remote areas with a limited number of users and restricted to low data-rate applications such as SMS. This situation will likely change with the introduction of voice and data services in the second half of 2025, or in the event of increased demand during large-scale outages of terrestrial infrastructure.
We observe that the correlation between RSRP and RSRQ is stronger in Starlink’s DS2D network than in the terrestrial network, with computed Pearson correlation coefficients of $0.66$ and $0.34$ respectively. This is consistent with the fact that network load is lower in the former than in the latter. With lower load, variations in the interference levels are smaller, which makes RSRP and RSRQ more correlated.
Lastly, the SINR measurements reveal a median value of 00 dB and an interdecile range of $16$ dB. These are again lower than the typical values observed in the terrestrial network, of $5$ dB and $20$ dB, respectively, as a result of the low-load, coverage-limited nature of the DS2D network.
Building on these observations, we next evaluate the potential performance of future DS2D data services.
IV-D Performance assessment of future DS2D services
To assess the performance of future DS2D services, we make use of the modified Shannon capacity formula with the measured SINR and available bandwidth, as described in Subsection III-C. The measured SINR values reported above are almost insensitive to the three options described for computing $m$, because large values of the SINR are infrequent in the DS2D network, and thus $m$ has a negligible impact on the results. Thus, the measured SINR mean and median values translate into spectral efficiency values of $0.79$ and $0.64$ bps/Hz, respectively. As the DS2D service currently operates in the PCS G block only, which offers $2\times 5$ MHz of bandwidth, this estimated spectral efficiency corresponds to an estimated downlink capacity of $4$ Mbps per beam, meaning the capacity a single user who occupies the entire beam bandwidth would experience outdoor in a remote or sparsely populated area.
While the estimated DS2D speed may be sufficient to support basic connectivity in remote or underserved areas, it remains significantly below the average download speed experienced in terrestrial networks. This gap underscores the complementary nature of DS2D services in its current phase, aligning with its intended role as a SCS solution. Nevertheless, performance is expected to improve in the coming months through a combination of increased radiated power, continued expansion of the satellite constellation, and greater bandwidth allocation per cell.
As noted above, the FCC authorized a $10$-dB increase in OOBE for DS2D transmissions in March 2025. Although measurements from April 2025 do not yet show any improvement in SINR or other metrics, this regulatory change could enable SINR gains of up to $10$ dB in the near future, provided that transmitter non-linearities do not necessitate an increased power back-off with the new OOBE limit, and that the increased radiated power does not lead to significantly higher interference, which would partially offset the SINR improvement. These assumptions are reasonable, as such effects are expected to be minimal in the current DS2D deployment, where transmit power is limited by regulatory limits and performance is constrained more by received power and noise than by interference. Assuming a full $10$-dB SINR gain, the estimated spectral efficiency rises to $2.35$ (mean) and $2.38$ bps/Hz (median)—a threefold increase which results in around $11.7$ Mbps. As expected, this aligns with the well-established principle that increasing SINR yields significant capacity gains in coverage-limited scenarios.
In Subsection III-C, a maximum MCS has been assumed corresponding to $256$-QAM. The 3GPP standard contemplates the use of even higher order modulation. However, this will not produce a noticeable capacity improvement. The reason is, again, the coverage-limited nature of the network: large values of SINR occur very infrequently with the existing SINR distribution, and thus the higher MCS are hardly ever used. This is the case even if the previously discussed $10$-dB increase is incorporated.
Throughput in the DS2D network may also be increased by expanding the available bandwidth. Assuming that the additional spectrum is paired with proportional transmit power, the spectral efficiency remains constant, and the per-beam capacity doubles. Starlink could pursue this by accessing more mobile spectrum (IMT) for the SCS services, but due to the risk of interference, this would require terrestrial operators to release spectrum, which is unlikely. Alternatively, Starlink may seek to use Mobile Satellite Service (MSS) spectrum and has, in fact, requested access to several bands around 2 GHz (specifically, the 1.6/2.4 GHz bands and the 2020–2025 MHz Earth-to-space band). However, this request was denied by the FCC in March 2024.444See FCC DA-24-300 In an additional $2\times 5$ MHz were granted in the future, the estimated mean throughput could increase from $4$ Mbps to $8$ Mbps and could potentially reach values close to $24$ Mbps when combined with gains from the recent increase in the OOBE limit.
In the longer term, satellite swarm array antennas show promising preliminary results in increasing the directivity of satellite beams [15], which allows for higher directivity gains and improved spatial multiplexing. Achieving this would require the launch and operation of many more satellites, which would also increase the overall throughput of DS2D networks and the effective capacity experienced by end users. Although the FCC denied Starlink’s request to authorize 30,000 DS2D-capable satellites, the company is still authorized to operate up to 7,500 such satellites, indicating substantial remaining potential for network densification.
V Conclusions
This paper presented the first empirical analysis of a commercially deployed DS2D network, focusing on Starlink’s Supplemental Coverage from Space Service in the U.S. in partnership with T-Mobile. Using large-scale crowdsourced measurements between October 2024 and April 2025, we characterized the spatio-temporal evolution and RAN performance of the DS2D network, which operated under an SMS-only configuration during the observed period.
Our results show a clear correlation between the number of DS2D-capable satellites launched and the number of unique cell identifiers observed, and spatial patterns that align well with the operator’s coverage gaps. We observed stable distributions of RSRP, RSRQ, and SINR across deployment phases, with signal power levels reflecting the expected limitations of satellite-based communication, and enhanced signal quality due to low traffic demand during this initial SMS-only phase. Based on SINR measurements, we estimate a current mean throughput of approximately $4$ Mbps, which is functional for basic services but significantly below terrestrial averages. Overall, the findings underscore the role of Starlink’s DS2D network as a solution for providing SCS.
Looking ahead, throughput and performance may improve to $24$ Mbps with higher radiated power (as recently authorized by the FCC), wider bandwidth allocation, and continued constellation expansion. However, these enhancements will depend on regulatory developments —particularly spectrum access— and on the evolution of service demand.
Finally, our work illustrates the value of crowdsourced data to monitor emerging network technologies. Beyond characterizing deployment and performance, such data can help detect underserved zones, assess coverage gaps, and inform infrastructure planning, offering a promising tool for both technical and policy research on next-generation mobile systems. Future work will involve monitoring the evolution of DS2D networks and services, tracking the evolution of Starlink’s network as new services are introduced, and exploring other emerging DS2D deployments.